# A posteriori closure of turbulence models: are symmetries preserved ?

### v1

@@ Abstract
Turbulence modeling remains a longstanding challenge in fluid dynamics. Recent advances in data-driven methods have led to a surge of novel approaches aimed at addressing this problem. This work builds upon [1], where we introduced a new closure for a shell model of turbulence using an a posteriori (or solver-in-the-loop) approach. Unlike most deep learning-based models, our method explicitly incorporates physical equations into the neural network framework, ensuring that the closure remains constrained by the underlying physics benefiting from enhanced stability and generalizability. In this paper, we further analyze the learned closure, probing its capabilities and limitations. In particular, we look at joint probability density functions to assess whether cross-correlations are well preserved or if just the mean behavior is captured. Additionally, we investigate the scale invariance of multipliers—ratios between adjacent shells—within the inertial range. Although our model excels in reproducing high-order observables such as flatness, it breaks this known symmetry near the cutoff, indicating a fundamental limitation. We discuss the implications of these findings for subgrid-scale modeling in 3D turbulence and outline directions for future research.
@@ 1 Introduction
Turbulence is ubiquitous in nature, from the coffee we stir in the morning to the primordial states of our universe [2]. Despite its ubiquity, understanding and modeling turbulence remains one of the most challenging tasks in classical physics. While we have known its governing equations for over a century, the full phenomenology of turbulence remains elusive. Simulating turbulence numerically is also very computationally heavy. The degrees of freedom scale as a power law of the Reynolds number, $Re=uL/\nu$, which means that studying very high Reynolds number turbulence remains impossible for the most part. One approach to reduce the degrees of freedom is Large Eddy Simulation (LES) [3], where a filter is applied, typically within the inertial range, to resolve only the largest, most energy-containing eddies. However, to close the LES system, the effects of unresolved subgrid scales must be modeled. This constitutes the LES modeling problem. This problem is often addressed phenomenologically by introducing an eddy diffusivity term to stabilize the system. These solutions, albeit useful for practical application, fail to capture the correct subgrid scale statistics. Moreover, although decades of research have been spent in LES modeling, there are still very fundamental open questions in the field. Can a subgrid-scale closure be formulated that accurately captures the statistics of the resolved scales, especially in high Reynolds number flows where intermittency plays a crucial role? In turbulence, velocity fields are not statistically self-similar due to intermittency, making it essential for closures to account for the multifractal nature of the subgrid scales.
Recently, deep learning (DL) approaches have gained attention for addressing the LES modeling problem. DL excels at approximating functions with unknown functional forms. By learning from data, DL methods can, albeit in a black-box manner, accurately approximate high-dimensional functions.
@@
To investigate the applicability of deep learning in this context, we use shell models of turbulence [4]. Shell models offer a simplified representation of 3D homogeneous isotropic turbulence (HIT), providing a tractable alternative to the full Navier-Stokes equations. Despite their simplicity, shell models preserve key turbulence phenomena, such as energy cascade, intermittency, anomalous scaling, and multifractal energy dissipation.
@@
Our primary focus is on the ability of subgrid-scale (SGS) models to accurately capture extreme events and rare fluctuations, which demand high precision, extensive statistical sampling, and long inertial ranges. Learning SGS models from data requires long temporal evolutions, presenting a fundamental challenge for LES of three-dimensional turbulence. This challenge is currently best addressed through shell models. Although LES of three-dimensional HIT remains the ideal test case, existing machine learning-based LES tools are hindered by their low resolution and the proximity of the cutoff to the forcing scales, preventing the emergence of anomalous behavior. Shell models, in contrast, provide a more suitable framework for rigorously testing the effectiveness of LES models under conditions of strong intermittency and for assessing whether solver-in-the-loop approaches can improve the tracking of extreme fluctuations over time. Furthermore, higher-order statistics—particularly of fourth order and beyond—naturally involve non-local interactions in Fourier space, implying that SGS fluctuations can influence the resolved scales in non-trivial ways. This observation is critical when aiming for accuracy beyond second-order moments, which are typically the focus of LES validation. Shell models thus offer a controlled environment to probe whether a closure model can correctly capture such high-order interactions and preserve the cross-scale dynamics.
@@
Several previous studies have explored closures within the shell model framework. Biferale et al. [5] initially approached the closure problem phenomenologically, using Kolmogorov multipliers. More recently, Ortali et al. [6] employed deep learning to learn the closure, while Domingues Lemos [7, 8] integrated probabilistic machine learning tools within the hidden symmetric framework of shell models [9] to address the closure. Building on these developments, we have also contributed to the field [1] through a solver-in-the-loop approach [10] for shell model closure based on the differentiable physics paradigm. This method unrolls the system in time during training, allowing the neural network to interact with a numerical solver for the governing equations over an arbitrary number of time steps before the loss function is computed. Rather than attempting to match the instantaneous, a priori evolution of the ground truth, we adopt an a posteriori approach, where the network learns how its errors propagate through time while receiving more realistic inputs. The induced closure was analyzed by examining high-order moments such as flatness up to the tenth order, demonstrating good agreement with the ground truth and unconditionally stable behavior.
This paper further extends this investigation, looking deeper into the learned closure. We will mainly focus on correlations between resolved and unresolved (i.e., predicted by the neural networks) variables, to assess whether the model accurately captures the cross-correlations, or whether it is only able to capture the mean behavior. Moreover, we will look into Kolmogorov multipliers statistics. These are ratios of velocity fluctuations between adjacent shells and have been shown to have universal statistics in the inertial range of fully developed turbulence. They were introduced by Kolmogorov, who hypothesized their self-similar statistics and scale invariance [11]. We will look at whether the scale invariance of these multipliers is broken by the subgrid-scale model, or whether the learned closure is able to preserve this symmetry, despite being trained purely on an $L^{2}$ norm of the velocity fields.
@@ 2 Shell models
Shell models are a class of turbulence models. They model homogeneous isotropic turbulence in Fourier space, considering only sparse coupling between the different shells. We discretize the 1D Fourier space with a geometric progression of shells that are equispaced on a logarithmic lattice, $k_{n}=k_{0}\lambda_{n}$, where one typically uses $\lambda\!=\!2$ and $k_{0}\!=\!1$. We consider the Sabra shell model [12], where the set of ordinary differential equations governing the dynamics of the different shells is given by:
@@
where $u_{n}\!\in\!\mathbb{C}$, represents the velocity fluctuations at a given shell $n$. The free parameters $a$, $b$, and $c$ are chosen such that $a+b+c=0$, which corresponds to conservation of energy $E\!=\!\sum_{n}|u_{n}|^{2}$ and helicity $H\!=\!\sum_{n}(a/c)^{n}|u_{n}|^{2}$ in the inviscid, unforced limit, mimicking the quadratic inviscid invariants in the Euler equations.
The first term in the RHS of Equation 1 represents the non-linear term, where we can see that only nearest neighbour and next-to-nearest neighbour interactions are preserved, making this model local in Fourier space; the second term represents the viscous term; and the third term represents large-scale forcing.
The left boundary conditions are $u_{-1}\!=\!u_{-1}\!=\!0$. For the fully resolved system, we integrate Equation 1 for shells $n\!=\!0,\ldots,N$, where $N$ is chosen with some margin of the dissipative cutoff scale, $N_{\eta}$, as to fully resolve all of the dynamics. In the case of closing the system in an LES fashion, we have to consider a cutoff shell $N_{c}$, after which we only have to model the shells $N_{c}\!+\!1$ and $N_{c}\!+\!2$ (the left BCs remain the same), as depicted in Figure 1. The velocity fields of these shells will appear in the equations for $u_{N_{c}-1}$ and $u_{N_{c}}$:
$\frac{du_{N_{c}}}{dt}$ $=i\Big{(}k_{N_{c}+1}u_{N_{c}+2}u_{N_{c}+1}^{*}-\frac{1}{2}k_{N_{c}}u_{N_{c}+1}u_{N_{c}-1}^{*}+\frac{1}{2}k_{N_{c}-1}u_{N_{c}-1}u_{N_{c}-2}\Big{)}-\nu k_{N_{c}}^{2}u_{N_{c}}$ (2a)
$\frac{du_{N_{c}-1}}{dt}$ $=i\Big{(}k_{N_{c}}u_{N_{c}+1}u_{N_{c}}^{*}-\frac{1}{2}k_{N_{c}-1}u_{N_{c}}u_{N_{c}-2}^{*}+\frac{1}{2}k_{N_{c}-2}u_{N_{c}-2}u_{N_{c}-3}\Big{)}-\nu k_{N_{c}-1}^{2}u_{N_{c}-1}$ (2b)
Figure 1: Energy spectrum $|u_{n}|^{2}$ as a function of the wavenumber $k_{n}$. The red-circled markers indicate the reduced system, while black crosses represent the fully resolved system. The inset zooms into the cutoff region, highlighting the nonlinear energy transfer through triadic interactions, with blue and green arrows indicating interactions for the two shells preceding the cutoff. It becomes clear that to close the system, only the two shells after the cutoff are needed to model. The x-axis is labeled in terms of $\log_{2}k_{n}$, with markers for the Kolmogorov scale $k_{\eta}$ and the maximum resolved wavenumber $k_{N}$.
@@ 3 A posteriori closure
To close the system, we use a deep neural network that predicts the two missing shells, $u_{N_{c}+1}$ and $u_{N_{c}+2}$, using as input the three preceding shells. Unlike an a priori approach, which trains the model to minimize one-step errors, we follow an a posteriori or solver-in-the-loop strategy, where the neural network operates within the evolving simulation. This allows it to learn not only accurate short-term predictions but also how its own errors propagate over time and influence the system’s long-term behavior.
A common challenge in machine learning is the distribution shift, i.e., when a model, after making several predictions, encounters states it has never been seen during training, leading to increasing errors. This is well known in autoregressive generative models, such as language models predicting text sequences [13]. Similarly, in physics-based modeling, an imperfect closure leads to accumulated deviations that push the system into untrained regions of phase space. By embedding the neural network directly within the governing equations, we ensure it learns to operate under realistic conditions, mitigating the effects of distribution shift. This approach leads to improved stability, better generalization, and a more physically consistent closure.
Let the reduced system $(\tilde{u}_{0},\tilde{u}_{1},\dots,\tilde{u}_{N_{c}})$ reside in the LES manifold $\mathcal{S}$, while the full system $(u_{0},u_{1},\ldots,u_{N})$ spans the reference manifold $\mathcal{R}$. The governing equations of the shell model define a dynamical system on $\mathcal{R}$, but large eddy simulations truncate the system to $\mathcal{S}$, introducing closure errors. To bridge this gap, we introduce a correction operator $\mathcal{C}:\mathcal{S}\!\rightarrow\!\bar{\mathcal{R}}$ parameterized by a neural network, where $\bar{\mathcal{R}}$ represents the truncated reference manifold, restricted to the resolved scales $(u_{0},u_{1},\ldots,u_{N_{c}})$. This operator predicts the unresolved shells $\tilde{u}_{N_{c}+1},\tilde{u}_{N_{c}+2}=\mathcal{C}(\tilde{u}_{N_{c}-2},\tilde{u}_{N_{c}-1},\tilde{u}_{N_{c}})$, enabling reconstruction of the full state in $\mathcal{C}$, via an explicit time integration of the missing terms: this corresponds to an additive correction to the resolved terms that are integrated through a fourth-order Runge-Kutta scheme.
@@
$\bm{h}^{(0)}=\bm{x},$
$\bm{h}^{(\ell)}=\sigma\left(\bm{W}^{(\ell)}\bm{h}^{(\ell-1)}+\bm{b}^{(\ell)}\right),\quad\text{for }\ell=1,\dots,L,$
$\tilde{\bm{u}}=\bm{W}^{(L+1)}\bm{h}^{(L)}+\bm{b}^{(L+1)}.$
Here, $\bm{W}^{(\ell)}$ and $\bm{b}^{(\ell)}$ are the weight matrix and bias vector for the $\ell$-th layer, while $\sigma(\cdot)$ is a nonlinear activation function, in our case ReLU. The final layer has a linear activation to ensure unrestricted output values.
@@
$\tilde{\bm{u}}(t+\Delta t)=\mathcal{P}(\tilde{\bm{u}}(t),\tilde{u}_{N_{c}+1},\tilde{u}_{N_{c}+2}),$
where $\tilde{\bm{u}}(t)=(\tilde{u}_{0},\dots,\tilde{u}_{N_{c}})_{t}$. Unlike a priori closure models trained to minimize single-step errors, our a posteriori approach integrates $\mathcal{C}$ recurrently into $\mathcal{P}$.
During training, the solver interacts with $\mathcal{C}$ over $m$ steps, exposing it to its own error dynamics. The loss function evaluates deviations from the reference trajectory across this horizon:
@@
where $\bm{u}^{\text{ref}}$ is obtained from a direct numerical simulation of $\mathcal{R}$. The optimization problem is then formulated as:
$\min_{\bm{\theta}}\mathcal{L}(\tilde{\bm{u}}^{\bm{\theta}},\bm{u}^{\text{ref}}),$
where $\bm{\theta}=\{\bm{W}^{(\ell)},\bm{b}^{(\ell)}\}_{\ell=1}^{L+1}$ represents the neural network parameters. This multi-step optimization ensures $\mathcal{C}$ learns to mitigate error accumulation and distribution shift, as illustrated in Figure 2.

### v2

@@ Abstract
Turbulence modeling remains a longstanding challenge in fluid dynamics. Recent advances in data-driven methods have led to a surge of novel approaches aimed at addressing this problem. This work builds upon our recent work [Phys. Rev. Fluids 10, 044602 (2025)], where we introduced a new closure for a shell model of turbulence using an a posteriori (or solver-in-the-loop) approach. Unlike most deep learning-based models, our method explicitly incorporates physical equations into the neural network framework, ensuring that the closure remains constrained by the underlying physics benefiting from enhanced stability and generalizability. In this paper, we further analyze the learned closure, probing its capabilities and limitations. In particular, we look at joint probability density functions between resolved and unresolved variables, as well as the scale invariance of multipliers (ratios between adjacent shells) within the inertial range. Although our model excels in reproducing high-order statistical moments, it breaks this known symmetry near the cutoff, indicating a fundamental limitation. We discuss the implications of these findings for subgrid-scale modeling in 3D turbulence and outline directions for future research.
@@ 1 Introduction
Turbulence is ubiquitous in nature, from the coffee we stir in the morning to the primordial states of our universe [17]. Despite its ubiquity, understanding and modeling turbulence remains one of the most challenging tasks in classical physics. While we have known its governing equations for over a century [10], the full phenomenology of turbulence remains elusive [43, 15]. Simulating turbulence numerically is also extremely computationally demanding [36, 21]. The degrees of freedom scale as a power law of the Reynolds number, $Re=UL/\nu$ (where $U$ is a characteristic velocity, $L$ the integral length scale, and $\nu$ the kinematic viscosity), which means that studying very high Reynolds number turbulence remains impossible for the most part. One approach to reduce the degrees of freedom is Large Eddy Simulation (LES) [35], where a filter is applied, typically within the inertial range, to resolve only the largest, most energy-containing eddies. However, to close the LES system, the effects of unresolved subgrid scales must be modeled. This constitutes the LES modeling problem. This problem is often addressed phenomenologically by introducing an eddy diffusivity term to stabilize the system [42]. This solution, albeit useful for practical application, fails to capture the correct subgrid scale statistics. Moreover, although decades of research have been spent in LES modeling [42, 30, 25, 29, 4, 13, 6], there are still fundamental open questions in the field. Can a subgrid-scale closure be formulated to accurately capture the statistics of the resolved scales, especially in high Reynolds number flows where intermittency [8] plays a crucial role? In turbulence, velocity fields are not statistically self-similar due to intermittency, making it essential for closures to account for the multifractal nature of the subgrid scales [40].
@@
Recently, deep learning (DL) [18] approaches have gained attention for addressing the LES modeling problem [13]. DL excels at approximating functions with unknown functional forms [9, 20]. By learning from data, DL methods can, albeit in a black-box manner [32], accurately approximate high-dimensional functions [1]. To investigate the applicability of DL in this context, we use shell models of turbulence [5] as our playground. Shell models offer a simplified representation of 3D homogeneous isotropic turbulence (HIT), providing a tractable alternative to the full Navier-Stokes equations. Despite their simplicity, shell models preserve key turbulence phenomena, such as energy cascade, intermittency, anomalous scaling, and multifractal energy dissipation.
@@
Our primary focus is on the ability of subgrid-scale (SGS) models to accurately capture extreme events and rare fluctuations, which demand high precision, extensive statistical sampling, and long inertial ranges. Learning SGS models from data requires long temporal evolutions, presenting a fundamental challenge for LES of three-dimensional turbulence. This challenge is currently best addressed through shell models. Although LES of three-dimensional HIT remains the ideal test case, existing machine learning-based LES tools are hindered by their low resolution and the proximity of the cutoff to the forcing scales, preventing the emergence of anomalous behavior. Shell models, in contrast, provide a more suitable framework for rigorously testing the effectiveness of LES models under conditions of strong intermittency and for assessing whether DL approaches can improve the tracking of extreme fluctuations over time. Furthermore, higher-order statistics–particularly of fourth order and beyond–naturally involve non-local interactions in Fourier space, implying that SGS fluctuations can influence the resolved scales in non-trivial ways. This observation is critical when aiming for accuracy beyond second-order moments, which are typically the focus of LES validation. Shell models thus offer a controlled environment to probe whether a closure model can correctly capture such high-order interactions.
@@
Several previous studies have explored closures within the shell model framework. Biferale et al. [4] initially approached the closure problem phenomenologically, using Kolmogorov multipliers. More recently, Ortali et al. [39] employed deep learning to learn the closure, while Domingues Lemos [12, 28] integrated probabilistic machine learning tools within the hidden symmetric framework of shell models [34] to address the closure. Building on these developments, we have also contributed to the field [16] through a solver-in-the-loop approach [44] based on the differentiable physics paradigm. This method unrolls the system in time during training, allowing the neural network to interact with a numerical solver for the governing equations over an arbitrary number of time steps before the loss function is computed. Rather than attempting to match the instantaneous, a priori evolution of the ground truth, we adopt an a posteriori approach, where the network learns how its errors propagate through time while receiving more realistic inputs. The induced closure was analyzed by examining high-order moments such as flatness up to the tenth order, demonstrating good agreement with the ground truth and unconditionally stable behavior.
This paper further extends our previous investigation by studying the learned closure in greater detail. While our earlier work [16] showed that the model reproduces the statistics of the resolved variables with high accuracy, here we test a stricter criterion: whether it can also recover nontrivial correlations between resolved and unresolved variables. To this end, we analyze joint probability density functions and other multi-scale statistics. We also investigate Kolmogorov multipliers [23], which are known to exhibit universal, scale-invariant statistics within the inertial range. These multipliers, especially those near the cutoff, intrinsically couple resolved and unresolved variables and therefore provide a sensitive diagnostic of whether the learned closure preserves or breaks this symmetry.
@@ 2 Shell models
Shell models are a class of turbulence models. They model homogeneous isotropic turbulence in Fourier space, considering only sparse coupling between the different shells. We discretize the 1D Fourier space with a geometric progression of shells that are equispaced on a logarithmic lattice, $k_{n}=k_{0}\lambda^{n}$, where one typically uses $\lambda\!=\!2$ and $k_{0}\!=\!1$. We consider the Sabra shell model [27], where the set of ordinary differential equations governing the dynamics of the different shells is given by:
@@
where $u_{n}\!\in\!\mathbb{C}$, represents the velocity fluctuations at a given shell $n$ and $^{*}$ denotes complex conjugation. We set the free parameters (a,b,c)=(1,1/2,1/2)(a,b,c)=(1,-1/2,-1/2) such that $a+b+c=0$, which corresponds to conservation of energy $E\!=\!\sum_{n}|u_{n}|^{2}$ and helicity $H\!=\!\sum_{n}(a/c)^{n}|u_{n}|^{2}$ in the inviscid, unforced limit.
The first term in the RHS of Equation 1 represents the non-linear term, where we can see that only nearest and next-to-nearest neighbour interactions are preserved, making this model local in Fourier space; the second term represents the viscous term; and the third term represents large-scale forcing. In our simulations, a constant forcing is applied on the first two shells.
The left boundary conditions (BCs) are $u_{-1}\!=\!u_{-2}\!=\!0$. For the fully resolved system, we integrate Equation 1 for shells $n\!=\!0,\ldots,N$, where $N$ is chosen with some margin of the dissipative cutoff scale, $N_{\eta}$, as to fully resolve all of the dynamics. In the case of closing the system in an LES fashion, we have to consider a cutoff shell $N_{c}$, after which we only have to model the shells $N_{c}\!+\!1$ and $N_{c}\!+\!2$ (the left BCs remain the same), as depicted in Figure 1. In all simulations, we set $N_{c}\!=\!14$, a cutoff chosen within the inertial range but safely above the dissipative cutoff. The velocity fields of the subgrid shells then appear explicitly in the equations for $u_{N_{c}}$ and $u_{N_{c}-1}$:
$\frac{du_{N_{c}}}{dt}$ $=i\Big(k_{N_{c}+1}u_{N_{c}+2}u_{N_{c}+1}^{*}-\frac{1}{2}k_{N_{c}}u_{N_{c}+1}u_{N_{c}-1}^{*}+\frac{1}{2}k_{N_{c}-1}u_{N_{c}-1}u_{N_{c}-2}\Big)-\nu k_{N_{c}}^{2}u_{N_{c}},$ (2a)
$\frac{du_{N_{c}-1}}{dt}$ $=i\Big(k_{N_{c}}u_{N_{c}+1}u_{N_{c}}^{*}-\frac{1}{2}k_{N_{c}-1}u_{N_{c}}u_{N_{c}-2}^{*}+\frac{1}{2}k_{N_{c}-2}u_{N_{c}-2}u_{N_{c}-3}\Big)-\nu k_{N_{c}-1}^{2}u_{N_{c}-1}.$ (2b)
Figure 1: Energy spectrum $|u_{n}|^{2}$ as a function of the wavenumber $k_{n}$. The red-circled markers indicate the reduced system, while black crosses represent the fully resolved system. The inset zooms into the cutoff region, highlighting the nonlinear energy transfer through triadic interactions, with blue and green arrows indicating interactions for the two shells preceding the cutoff. It becomes clear that to close the system, only the two shells after the cutoff are needed to model. The x-axis is labeled in $\log_{2}k_{n}$, along with the Kolmogorov scale $k_{\eta}$ and maximum resolved wavenumber $k_{N}$.
@@ 3 A posteriori closure
To close the system, we use a deep neural network that predicts the two missing shells, $u_{N_{c}+1}$ and $u_{N_{c}+2}$, using as input the three preceding shells (enough to fix the flux locally). Unlike an a priori approach, which trains the model to match instantaneously the subgrid terms, we follow an a posteriori or solver-in-the-loop strategy, where the neural network operates within the evolving simulation. This allows it to learn not only accurate short-term predictions but also how its own errors propagate over time and influence the system’s long-term behavior.
A key motivation for adopting an a posteriori strategy is the presence of distribution shift [41], that is, a mismatch between the statistical distributions encountered during training and deployment. Here, we denote by $x$ the model inputs and by $y$ the corresponding outputs or targets. Three main forms are typically distinguished. (i) Covariate shift, where the input distribution changes, $p_{\text{train}}(x)\neq p_{\text{test}}(x)$, while the conditional mapping remains the same, $p(y|x)$; (ii) Prior or label shift, where the output distribution differs, $p_{\text{train}}(y)\neq p_{\text{test}}(y)$, but $p(x|y)$ is preserved; and (iii) Concept shift, where the underlying relationship between inputs and outputs itself changes, $p_{\text{train}}(y|x)\neq p_{\text{test}}(y|x)$.
In our setting, the shift arises as a dynamically induced covariate shift. During solver deployment, small prediction errors in the learned closure accumulate and perturb the system dynamics, thereby altering the distribution of the model inputs—i.e., the resolved variables provided to the closure at each time step. Although the physical relation governing the coupling between resolved and unresolved scales remains invariant, the model gradually encounters regions of phase space that were rare or unseen during training. This phenomenon parallels what is observed in autoregressive generative models, such as language models, where compounding errors cause trajectories to drift away from the data manifold [19]. Embedding the neural network directly within the governing equations during training (the solver-in-the-loop strategy) exposes it to its own prediction-induced state distribution, thus reducing the mismatch between training and deployment distributions. This mitigates the covariate shift, leading to improved stability, generalization, and a more physically consistent closure.
To formalize this approach, let the reduced system $(\tilde{u}_{0},\tilde{u}_{1},\dots,\tilde{u}_{N_{c}})$ reside in the LES manifold $\mathcal{S}$, while the full system $(u_{0},u_{1},\ldots,u_{N})$ spans the reference manifold $\mathcal{R}$. The governing equations of the shell model define a dynamical system on $\mathcal{R}$, but large eddy simulations truncate the system to $\mathcal{S}$, introducing closure errors. To bridge this gap, we introduce a correction operator $\mathcal{C}:\mathcal{S}\!\rightarrow\!\bar{\mathcal{R}}$ parameterized by a neural network, where $\bar{\mathcal{R}}$ represents the truncated reference manifold, restricted to the resolved scales $(u_{0},u_{1},\ldots,u_{N_{c}})$. This operator predicts the unresolved shells $\tilde{u}_{N_{c}+1},\tilde{u}_{N_{c}+2}=\mathcal{C}(\tilde{u}_{N_{c}-2},\tilde{u}_{N_{c}-1},\tilde{u}_{N_{c}})$, enabling reconstruction of the full state in $\mathcal{C}$, via an explicit time integration of the missing terms: this corresponds to an additive correction to the resolved terms that are integrated through a fourth-order Runge-Kutta scheme [26].
@@
$\bm{h}^{(0)}$ $=\bm{x},$ (3a)
$\bm{h}^{(\ell)}$ $=\sigma\!\left(\bm{W}^{(\ell)}\bm{h}^{(\ell-1)}+\bm{b}^{(\ell)}\right),\quad\text{for }\ell=1,\dots,L,$ (3b)
$\tilde{\bm{u}}$ $=\bm{W}^{(L+1)}\bm{h}^{(L)}+\bm{b}^{(L+1)}.$ (3c)
Here, $\bm{W}^{(\ell)}$ and $\bm{b}^{(\ell)}$ are the weight matrix and bias vector for the $\ell$-th layer, while $\sigma(\cdot)$ is a nonlinear activation function, in our case ReLU [38]. The final layer has a linear activation to ensure unrestricted output values.
@@
$\tilde{\bm{u}}(t+\Delta t)=\mathcal{P}(\tilde{\bm{u}}(t),\tilde{u}_{N_{c}+1},\tilde{u}_{N_{c}+2}),$ (4)
where $\tilde{\bm{u}}(t)=(\tilde{u}_{0},\dots,\tilde{u}_{N_{c}})|_{t}$. Unlike a priori closure models trained to minimize instantaneous errors, our a posteriori approach integrates $\mathcal{C}$ recurrently into $\mathcal{P}$. During training, the solver interacts with $\mathcal{C}$ over $m$ steps, exposing it to its own error dynamics. The loss function evaluates deviations from the reference trajectory across this horizon:
@@
where $\bm{u}^{\text{ref}}$ is obtained from a direct numerical simulation of $\mathcal{R}$ and $\tilde{\bm{u}}^{\bm{\theta}}$ is the LES solution with the neural network parameters $\bm{\theta}=\{\bm{W}^{(\ell)},\bm{b}^{(\ell)}\}_{\ell=1}^{L+1}$. The optimization problem is then formulated as:
$\min_{\bm{\theta}}\mathcal{L}(\tilde{\bm{u}}^{\bm{\theta}},\bm{u}^{\text{ref}}).$
This multi-step optimization ensures $\mathcal{C}$ learns to mitigate error accumulation and distribution shift, as illustrated in Figure 2.
@@
The MLP used for the closure has seven hidden layers of 256 neurons each. Training is performed with the Adam optimiser [22], using a learning rate of $10^{-4}$. A detailed summary of the numerical and training parameters as well as the algorithm for training is provided in Appendix A. For more detailed information regarding the training procedure we refer the readers to Ref. [16].